OpenStudy (anonymous):

Simplify the following expression : 6^3/6^-1

OpenStudy (anonymous):

6^3/6^-1= 6^3*6^1=6^(3+1)=6^4 ok?

OpenStudy (anonymous):

The choices are: 1/36 1/1296 1296 36

OpenStudy (anonymous):

6*6*6*6

OpenStudy (anonymous):

ok ? @Sophia.13

OpenStudy (anonymous):

Yeah, thanks.

OpenStudy (anonymous):

(: